# On the $\Gamma$-limit of weighted fractional energies

### v1

@@ 4.1. Stability of Hilbertian gradient flows
Let $\mathscr{H}$ be a Hilbert space endowed with scalar product $\langle\cdot,\cdot\rangle_{\mathscr{H}}$ and norm $\|\cdot\|_{\mathscr{H}}$. Given $\mathscr{F}\colon\mathscr{H}\to(-\infty,+\infty]$, we let $\mathscr{D}(\mathscr{F})=\{\,x\in\mathscr{H}\colon\mathscr{F}(x)<+\infty\,\}$ and
@@
We recall the following result, which is a particular case of [CdLKNP23]*Prop. 3.7. Here and below, given any vector space $\mathscr{V}$, we let $\mathscr{V}^{*}$ be the algebraic dual space of $\mathscr{V}$, and we let $\mathscr{V}^{\prime}$ be the topological dual space of $\mathscr{V}$.
@@
Let $\mathscr{K}$ be a dense subspace of $\mathscr{H}$. Let $\mathscr{F}\colon\mathscr{H}\to(-\infty,+\infty]$ be a proper, convex, and strongly lower semicontinuous functional and let $x\in\mathscr{D}(\mathscr{F})$. If there exists $T\in\mathscr{K}^{*}$ such that
@@
Let $\mathscr{F}_{n}\colon\mathscr{H}\to(-\infty,+\infty]$ be a proper, strongly lower semicontinuous, convex and positive functional for every $n\in\mathbb{N}$. Assume the following:
@@
    $(\mathscr{F}_{n})_{n\in\mathbb{N}}$ $\Gamma$-converges to some proper functional $\mathscr{F}_{\infty}\colon\mathscr{H}\to(-\infty,+\infty]$ with respect to the strong $\mathscr{H}$-convergence;
@@
    admits a unique solution $x_{n}\colon[0,T]\to\mathscr{H}$ for every $n\in\mathbb{N}\cup\{\infty\}$;
@@
4.2. Proof of theorem 1.2
[…]
@@
    if $[\tilde{u}]_{s,2,f}<\infty$ for some $s\in(0,1)$, then
@@
We are now ready to prove of theorem 1.2.
Proof of theorem 1.2.
Since $f_{n}\to f$ uniformly and $f>0$, the functionals $\mathscr{F}_{n}(u)=(1-s_{n})[\tilde{u}]_{s_{n},2,f_{n}}^{2}$ are positive for $n\gg 1$. Further, they are easily shown to be convex. By theorem 1.1(ii)(iii), $(\mathscr{F}_{n})_{n\in\mathbb{N}}$ $\Gamma$-converges to the functional $\mathscr{F}(u)=K_{d,2}\|\nabla u\|^{2}_{2,f^{2}}$ in the strong topology of $L^{2}(\Omega)$, whereas by theorem 1.1(i) every sequence $(u_{n})_{n\in\mathbb{N}}\subset L^{2}(\Omega)$ such that $\sup_{n\in\mathbb{N}}\mathscr{F}_{n}(u_{n})<\infty$ admits a strong $L^{2}$-limit $u\in L^{2}(\Omega)$. The conclusion hence follows by theorems 4.2 and 4.1 combined with proposition 4.3. ∎

### v2

@@ 4.1. Stability of Hilbertian gradient flows
Let $\mathscr{H}$ be a Hilbert space endowed with scalar product $\langle\cdot,\cdot\rangle_{\mathscr{H}}$ and norm $\|\cdot\|_{\mathscr{H}}$. Given :($\mathscr{F}\colon\mathscr{H}\to$(-$\infty$$\infty$], we let $\mathscr{D}(\mathscr{F})=\{\,x\in\mathscr{H}\colon\mathscr{F}(x)<+\infty\,\}$ and
@@
We recall the following result, which is a particular case of [CdLKNP23]*Prop. 3.7. Here and below, given any vector space $\mathscr{V}$, we let $\mathscr{V}^{*}$ be the algebraic dual space of $\mathscr{V}$ and $\mathscr{V}^{\prime}$ the topological dual space of $\mathscr{V}$.
@@
Let $\mathscr{K}$ be a dense subspace of $\mathscr{H}$. Let :($\mathscr{F}\colon\mathscr{H}\to$(-$\infty$$\infty$] be a proper, convex, and strongly lower semicontinuous functional and let $x\in\mathscr{D}(\mathscr{F})$. If there exists $T\in\mathscr{K}^{*}$ such that
@@
[…]
@@
    $(\mathscr{F}_{n})_{n\in\mathbb{N}}$ $\Gamma$-converges to some proper functional :($\mathscr{F}_{\infty}\colon\mathscr{H}\to$(-$\infty$$\infty$] with respect to the strong $\mathscr{H}$-convergence;
@@
    admits a unique solution $x_{n}\in H^{1}([0,T];\mathscr{H})$ for every $n\in\mathbb{N}\cup\{\infty\}$;
@@
4.2. Proof of Theorem˜1.2
The validity of Theorem˜1.2 follows by combining the abstract results above with the following proposition. Here and below, $\langle\cdot,\cdot\rangle$ denotes the standard scalar product in $L^{2}(\mathbb{R}^{d})$.
@@
    if $\tilde{u}]_{s,2,f}<\infty$ for some $s\in(0,1)$, then
@@
We are now ready to prove Theorem˜1.2.
Proof of Theorem˜1.2.
Since $f_{n}\to f$ uniformly and $f>0$, the functionals given by $\mathscr{F}_{n}(u)=(1-s_{n})[\tilde{u}]_{s_{n},2,f_{n}}^{2}$ are positive for $n\gg 1$. Further, they are easily shown to be convex. By Theorem˜1.1(ii)(iii), $(\mathscr{F}_{n})_{n\in\mathbb{N}}$ $\Gamma$-converges to the functional $\mathscr{F}(u)=K_{d,2}\|\nabla u\|^{2}_{2,f^{2}}$ in the strong topology of $L^{2}(\Omega)$, whereas by Theorem˜1.1(i) every sequence $(u_{n})_{n\in\mathbb{N}}\subset L^{2}(\Omega)$ such that $\sup_{n\in\mathbb{N}}\mathscr{F}_{n}(u_{n})<\infty$ admits a strong $L^{2}$-limit $u\in L^{2}(\Omega)$. The conclusion hence follows by Theorems˜4.2 and 4.1 combined with Proposition˜4.3. ∎